OpenStudy (anonymous):

Simplify 1+cos4theta using double angle formula

OpenStudy (anonymous):

2*(cos{2theta})^2

OpenStudy (anonymous):

I dont follow, where as the 1 gone? and which formula did you use?

OpenStudy (anonymous):

\[2\cos^2\theta =\cos2\theta +1\]

OpenStudy (anonymous):

i think you understand